# ACE-RTL: When Agentic Context Evolution Meets RTL-Specialized LLMs

### v1

@@ 4.5. Results on VerilogEval
Table 2. Pass rate results on VerilogEval-Human-v2 — ACE-RTL-G denotes the standalone ACE-RTL-Generator model.
@@
CraftRTL Claude4-Sonnet ACE-RTL-G ScaleRTL VerilogCoder ACE-RTL
$68.0$ $73.0$ $\mathbf{73.8}$ $93.6$ $94.2$ $\mathbf{95.5}$
For completeness, we also evaluate our approach on VerilogEval-Human-v2. To ensure fair and rigorous comparison, we use Pass@1 for standalone models and APR for agentic methods. As shown in Table 2, ACE-RTL-G outperforms both the strong RTL model CraftRTL and the frontier model Claude4-Sonnet, demonstrating the strength of our RTL-specialized generator. Moreover, ACE-RTL achieves the highest pass rate among all agentic baselines, surpassing ScaleRTL and VerilogCoder. These results show that ACE-RTL is not overfitted to CVDP but instead captures transferable RTL reasoning and coding capabilities, maintaining strong performance on earlier RTL benchmarks.
5. Conclusion
This work introduces ACE-RTL, a framework that integrates an RTL-specialized model with ACE for accurate RTL code generation. ACE-RTL achieves SoTA results on multiple RTL benchmarks, surpassing both frontier proprietary and RTL-specialized models.

### v2

@@ IV-F Results on VerilogEval
Table 3: Pass rate results on VerilogEval-Human-v2 — ACE-RTL-G denotes the standalone ACE-RTL-Generator model.
@@
CraftRTL Claude4 ACE-RTL-G VerilogCoder MAGE ACE-RTL
$68.0$ $73.0$ $\mathbf{73.8}$ $94.2$ $\mathbf{95.5}$ $\mathbf{95.5}$
For completeness, we also evaluate our approach on VerilogEval-Human-v2. We use Pass@1 for standalone models and APR for agentic methods. As shown in Table 3, ACE-RTL-G outperforms prior RTL LLMs such as OriGen and CraftRTL. Moreover, ACE-RTL achieves the highest pass rate among all agentic baselines. These results show that ACE-RTL is not overfitted to CVDP but instead captures transferable RTL reasoning and coding capabilities, maintaining strong performance on earlier RTL benchmarks.
V Failure Analysis and Future Work
Although ACE-RTL achieves strong results on CVDP, it does not yet reach a $100\%$ pass rate. We observe two main sources of remaining failures. First, some datapoints require fine-grained temporal diagnosis, where textual simulation feedback alone is insufficient to localize the true timing-related bug. Second, some failures arise from ambiguity in the specification, or from subtle mismatches between the intended design behavior and the behavior emphasized by the test cases. These cases require not only debugging against observed failures, but also reasoning about the true design intent described in the specification.
One promising direction for addressing these failures is to incorporate waveform information, which can reveal temporal mismatches more directly than textual logs. However, we intentionally exclude waveform information in this work to evaluate whether ACE-RTL can solve RTL tasks primarily through RTL knowledge acquired via domain-specific training and specification-grounded reasoning, using only limited non-waveform simulation feedback. Although waveform-driven debugging often boosts the pass rate of LLM agents, it also risks overfitting to observed test behavior while underutilizing the specification itself, especially in cases with specification ambiguity or specification-test mismatch. We therefore view waveform information not as a substitute for specification-grounded reasoning, but as a potentially useful auxiliary signal that should be integrated into ACE-RTL in a careful and principled way to avoid overfitting to test cases. We leave such integration as an important future direction.
VI Conclusion
This work introduces ACE-RTL, a framework that integrates an RTL model with ACE and achieves SoTA results on multiple RTL benchmarks. An interesting future direction is to automate testbench generation and explore waveform-enhanced agentic reasoning.